OpenStudy (anonymous):

if h(x)=2g(x) - (1)/(g(x)) =2 and g'(0)=-1, find the exact value of h'(0). please please please help and show all steps!

OpenStudy (jamesj):

Suppose h(x) = 1/g(x). What is h'(x) equal to?

OpenStudy (jamesj):

By the chain rule, h' = dh/dg . dg/dx = -1/g^2 . g' i.e., h'(x) = -g'(x)/g(x)^2 Therefore h'(0) = -g'(0)/g(0)^2 Now apply this procedure to your function h(x)